# Do Current Language Models Support Code Intelligence for R Programming Language?

### v2

@@ 5.3. Intra- and Cross-Project
Analyzing the effects of intra- and cross-project contexts on model’s performance for R language is complex and nuanced. This aspect of model training presents unique challenges and opportunities that can significantly impact the effectiveness of pre-trained models and LLMs working with R.
Cross-project context presents an greater challenge. R is used across various domains, from bioinformatics to finance, each with its own specialized libraries, data structures, and best practices. A model trained on a broad spectrum of R projects might produce generalized code that fails to account for domain-specific requirements or optimizations. For example, code generated for a bioinformatics project might not inherently understand the need for efficient large-scale data processing that is crucial in genomics research. Another significant challenge lies in understanding and maintaining project-specific logic and business rules. Many R projects in professional settings implement complex analytical pipelines or decision-making processes that are unique to their use case. These often rely on domain knowledge that is not explicitly coded but is crucial for correct implementation.
@@ 6. Implications and Future Directions
Furthermore, researchers and practitioners can devise strategies to enhance model performance, specifically with the lack of huge datasets. This could also include optimizing the selection process and adaptation of Code-PLMs or LLMs for different tasks in R. As observed in RQ5, the LLMs performance on code generation for R is falling behind Python. Specific techniques could be developed to generate acceptable code in R. The code generation and creating a benchmark in Tidy-verse style is also another area of research.
@@
Second, our study provides valuable guidance for R language users who seek to leverage the current Code-PLMs. The comparison and evaluation of different pre-trained models in the R environment shed light on their relative strengths and weaknesses. This information empowers practitioners to make informed decisions when selecting the most suitable model for their specific use cases, considering factors such as model accuracy and compatibility with R programming styles. The practitioner and resercher should be aware of the limitations and strengths of the models when intend to use them for R.
@@
As discussed earlier, the Base and Tidy styles have different syntax and combining them in the training dataset could lower the performance. Careful curation of training data to ensure balanced representation of both Tidy-verse and Base R code is crucial, even separating them for specific tasks. Additionally, models might benefit from explicit tagging or context-awareness to distinguish between Tidy-verse and Base R code segments. This could enable more accurate and context-appropriate code generation and completion. Ultimately, the goal would be to develop models that can seamlessly work with both paradigms, understanding their strengths and appropriately suggesting or generating code that aligns with the user’s preferred style or the specific requirements of the task at hand.
@@
Another implication for researchers is exploring ways to develop datasets and techniques to improve the performance of models for project-specific or R specific domains. One potential solution is to develop models that can intake and process broader project context, e.g., by analyzing entire codebases or project documentation before making suggestions or generating code. This could help the model align its outputs more closely with project-specific conventions and requirements. Another approach might involve creating domain-specific models or fine-tuning general R models for particular fields of application. This could help in generating more relevant and optimized code for specific types of R projects, whether they are in finance, healthcare, or scientific research.
@@
The integration of pre-trained models in R opens up opportunities for their application for other low-resource (i.e., the amount of available data is limited) programming languages, and software engineering tasks. Researchers and practitioners working on these topics can draw upon our findings to explore and exploit the potential of pre-trained models within the R ecosystem, leading to improved solutions and advancements in their respective areas, for other code intelligent tasks and other under-studied programming languages, including scientific software.
@@ 7. Threats to Validity
Internal validity: The internal validity in our work could be related to the dataset collection process and pre-processing steps. In our study, the quality of data is a significant concern. The code and associated documentation within R packages can vary widely in quality. Some packages may be meticulously maintained with high-quality code and comprehensive documentation, providing valuable insights for model training. However, others may be less well-maintained, with misleading, inaccurate, or sparse documentation. This inconsistency can lead to a model that is trained on a mixture of high and low-quality data, potentially reducing its effectiveness and accuracy. To mitigate this issue, it is crucial to implement robust data cleaning and preprocessing steps, possibly including manual review and curation of the dataset. To achieve this, we consulted constantly with an R expert, who has more than ten years of research and development experience in R, ensuring the quality of the collected dataset and the soundness of the approach, specifically the decision on Base or Tidy-verse R repositories. We also manually evaluated the collected data at different steps, to ensure our developed process leads to the desired data, e.g., the code and comments are tokenized correctly. For the tokenization part, we used several tools, and each failed in one part. Finally, we found a reliable grammar rule for R that could be used with a Tree-sitter. The obtained dataset was investigated by the first author and another student, ensuring that the code and comments were separated correctly. These steps were taken in addition to following the advice of the R expert and previously published works to select the inclusion and exclusion criteria to scrape the related GitHub repositories.
@@
Construct validity: We use the commonly used evaluation metrics for code summarization and method name prediction. For code generation we also used a benchmark dataset and the Pass@K metrics, which are well studied in the related literature. These metrics are well known in NLP and SE research and therefore we do not anticipate such validity threats.
@@ 8. Conclusion and Future Studies
In this study, we explored the capabilities of the Code-PLMs for the two tasks of code summarization and method name prediction for R. This is the first attempt towards code intelligence in R for the mentioned tasks. We also examined LLMs for code generation task for R programming language. Both LLMs and Code-PLMs consistently experience varying degrees of performance degradation when processing R code compared to other programming languages. The challenge stems in different syntax of the R styles as well as its project-specific and package-based nature. Interestingly, increasing the dataset size or adding other programming languages for multi-lingual fine-tuning not necessarily helps in improving the performance.
@@ Acknowledgment
This research is support by a grant from Natural Sciences and Engineering Research Council of Canada RGPIN-2019-05175.

### v3

@@ 5.3. Intra- and Cross-Project
Analyzing the effects of intra- and cross-project contexts on a model’s performance for the R language is complex and nuanced. This aspect of model training presents unique challenges and opportunities that can significantly impact the effectiveness of pre-trained models and LLMs working with R.
Cross-project context presents a greater challenge. R is used across various domains, from bioinformatics to finance, each with its own specialized libraries, data structures, and best practices. A model trained on a broad spectrum of R projects might produce generalized code that fails to account for domain-specific requirements or optimizations. For example, code generated for a bioinformatics project might not inherently understand the need for efficient large-scale data processing that is crucial in genomics research. Another significant challenge lies in understanding and maintaining project-specific logic and business rules. Many R projects in professional settings implement complex analytical pipelines or decision-making processes that are unique to their use case. These often rely on domain knowledge that is not explicitly coded but is crucial for correct implementation.
5.4. Discussion on Model Behavior
All of our baseline models (CodeBERT, GraphCodeBERT, UniXCoder, and CodeT5) differ in how they learn and leverage code representations. First, each model’s pretraining objective shapes what it emphasizes: BERT-like objectives (as in CodeBERT and GraphCodeBERT) encourage strong token-level encoding, while T5-style objectives (as in CodeT5) promote sequence-to-sequence generation, naturally aligning with summarization tasks. Second, language-specific syntax affects performance: more verbose or structurally consistent languages (e.g., Java) often suit models that rely on explicit tokens and clear code patterns, whereas dynamic or less standard languages (e.g., Ruby, or different R styles) can pose unique challenges. Third, internal attention patterns vary by architecture: GraphCodeBERT’s additional graph representation may highlight data flows, whereas CodeT5’s text-to-text design focuses on coherent generation. Finally, fine-tuning data coverage can cause certain models to overfit to common constructs or underperform on outlier patterns. Together, these factors may explain performance fluctuation across different programming languages. However, a comprehensive study separate from this work is required to delve into the exact causes and their effects.
@@ 6. Implications and Future Directions
Furthermore, researchers and practitioners can devise strategies to enhance model performance, specifically with the lack of huge datasets. This could also include optimizing the selection process and adaptation of Code-PLMs or LLMs for different tasks in R. As observed in RQ5, the LLMs’ performance on code generation for R is falling behind Python. Specific techniques could be developed to generate acceptable code in R. The code generation and creating a benchmark in Tidy-verse style is also another area of research.
@@
Second, our study provides valuable guidance for R language users who seek to leverage the current Code-PLMs. The comparison and evaluation of different pre-trained models in the R environment shed light on their relative strengths and weaknesses. This information empowers practitioners to make informed decisions when selecting the most suitable model for their specific use cases, considering factors such as model accuracy and compatibility with R programming styles. The practitioner and researcher should be aware of the limitations and strengths of the models when intend to use them for R.
@@
As discussed earlier, the Base and Tidy styles have different syntax, and combining them in the training dataset could lower the performance. Careful curation of training data to ensure balanced representation of both Tidy-verse and Base R code is crucial, even separating them for specific tasks. Additionally, models might benefit from explicit tagging or context-awareness to distinguish between Tidy-verse and Base R code segments. This could enable more accurate and context-appropriate code generation and completion. Ultimately, the goal would be to develop models that can seamlessly work with both paradigms, understanding their strengths and appropriately suggesting or generating code that aligns with the user’s preferred style or the specific requirements of the task at hand.
@@
Another implication for researchers is exploring ways to develop datasets and techniques to improve the performance of models for project-specific or R specific domains. One potential solution is to develop models that can intake and process broader project contexts, e.g., by analyzing entire codebases or project documentation before making suggestions or generating code. This could help the model align its outputs more closely with project-specific conventions and requirements. Another approach might involve creating domain-specific models or fine-tuning general R models for particular fields of application. This could help in generating more relevant and optimized code for specific types of R projects, whether they are in finance, healthcare, or scientific research.
@@
The integration of pre-trained models in R opens up opportunities for their application in other low-resource (i.e., the amount of available data is limited) programming languages and software engineering tasks. Researchers and practitioners working on these topics can draw upon our findings to explore and exploit the potential of pre-trained models within the R ecosystem, leading to improved solutions and advancements in their respective areas, for other code-intelligent tasks and other under-studied programming languages, including scientific software.
@@ 7. Threats to Validity
An additional threat to validity arises from our exclusive reliance on GitHub repositories. While GitHub was selected to facilitate direct identification of Roxygen2 parser usage and to streamline quality-focused filtering, depending solely on GitHub may overlook specialized R usage scenarios, particularly in fields like bioinformatics or finance, where code is often shared on CRAN or Bioconductor. As a result, our dataset might not fully capture these domain-specific practices.
Internal validity: The internal validity in our work could be related to the dataset collection process and pre-processing steps. In our study, the quality of data is a significant concern. The code and associated documentation within R packages can vary widely in quality. Some packages may be meticulously maintained with high-quality code and comprehensive documentation, providing valuable insights for model training. However, others may be less well-maintained, with misleading, inaccurate, or sparse documentation. This inconsistency can lead to a model that is trained on a mixture of high and low-quality data, potentially reducing its effectiveness and accuracy. To mitigate this issue, it is crucial to implement robust data cleaning and preprocessing steps, possibly including manual review and curation of the dataset.
To achieve this, we consulted constantly with an R expert, who has more than ten years of research and development experience in R, ensuring the quality of the collected dataset and the soundness of the approach, specifically the decision on Base or Tidy-verse R repositories. We also manually evaluated the collected data at different steps to ensure our developed process leads to the desired data, e.g., the code and comments are tokenized correctly.
For the tokenization part, we used several tools, and each failed in one part. Finally, we found a reliable grammar rule for R that could be used with a Tree-sitter. The obtained dataset was investigated by the first author and another student, ensuring that the code and comments were separated correctly. These steps were taken in addition to following the advice of the R expert and previously published works to select the inclusion and exclusion criteria to scrape the related GitHub repositories.
@@
Construct validity: We use the commonly used evaluation metrics for code summarization and method name prediction. For code generation we also used a benchmark dataset and the Pass@K metrics, which are well studied in the related literature. These metrics are well known in NLP and SE research and therefore, we do not anticipate such validity threats.
@@ 8. Conclusion and Future Studies
In this study, we explored the capabilities of the Code-PLMs for the two tasks of code summarization and method name prediction for R. This is the first attempt towards code intelligence in R for the mentioned tasks. We also examined LLMs for the code generation task for the R programming language. Both LLMs and Code-PLMs consistently experience varying degrees of performance degradation when processing R code compared to other programming languages. The challenge stems from the different syntax of the R styles as well as its project-specific and package-based nature. Interestingly, increasing the dataset size or adding other programming languages for multi-lingual fine-tuning does not necessarily help in improving the performance.
@@ Acknowledgment
This research is supported by a grant from Natural Sciences and Engineering Research Council of Canada RGPIN-2019-05175.